# The Galactic population of canonical pulsars II

### v1

@@ 3.1.2 With the death line
Adding a death line decreases the number of pulsars detected in radio compared to the situation without a death line, see Table 7.
@@
Figure 7: Density plot of the $P-\dot{P}$ diagram in comparison with observations
The death line improves the fit as seen in Fig. 5 for the distribution of spin period and compared with Fig. 2, even though there is a slight excess of pulsars which have spin period between $10^{-0.5}$ s and $1$ s. In Fig 6, these pulsars in excess lie at $\dot{P}$ between $10^{-18}$ and $10^{-17}$, where there is no data and which correspond to the oldest pulsars in the simulation. Therefore the oldest pulsars simulated here do not correspond to any real data.
In addition, there are too many pulsars close to the death line compared to the observations, there is a pile up effect created by the death line which is not in the observations.
The goodness-of-fit can be checked using a density plot of the $P-\dot{P}$ diagram as in Johnston & Karastergiou (2017) instead of putting points for individual pulsars. In order to construct such diagrams we binned the data evenly in a log scale in $P$ and $\dot{P}$ and count the number of pulsars in each bin. This leads to a 2D histogram to be compared with the observed density plot 2D histogram. Because the total number of simulated pulsars $N_{totsim}$ is different from the total number of observed pulsars $N_{totobs}$, we normalise the number of pulsars in each bin in order to compare proportions. We therefore introduce the quantity $R$ as follows
@@
With this definition the number $R$ remains in the interval $[-1,+1]$. $N_{sim}$ is the number of pulsar count in a bin of the simulation, $N_{obs}$ the number of pulsar detected in a bin of the observations. A good fit in a bin correspond to a value of $R$ close to 00, meaning that the simulation results are close to the observations. Inspecting fig. 7, we note a stronger deviation from the observations at the boundaries of the density plot, for instance it confirms too many pulsars are piling up close to the death line. The high values of $R$ at these boundaries originates from the weak number of pulsars counted in this bins leading to bad statistics and large variations in the proportions, a side effect of the density plot approach. The fact that the colors are very dark, shows that the proportion of pulsars in most of the bins are not good.
A Kolmogorov Smirnov (KS) test (Smirnov, 1948) was conducted on both $P$ and $\dot{P}$ which gave p-values substantially below 0.05 respectively when no death line is implemented. While we obtained p-values of 0.05 and another one well below 0.05 respectively for $P$ and $\dot{P}$ with a death line implementation. The null hypothesis that observations and simulations originate from the same distribution for the spin period can not be rejected (p-value $\geq$ 0.05) only when a death line is considered, but is clearly rejected for $\dot{P}$ in both case.
@@ 3.2.1 With the death line
The parameters used for these runs are shown in Table 5. This new set of simulations is intended to verify if generating pulsars younger than $4.1\times 10^{7}$ yr is sufficient, since we observe a small excess of very old pulsars. The death line is also taken into account.
@@
Table 8 shows that a few less pulsars were detected in this simulation compared to the one with the death line and 10 millions pulsars (see Table 7). Moreover, inspecting the histograms for spin period and spin period derivative respectively in Fig. 8 and 9, we conclude that observations and simulations are very similar in both histograms. A KS test was conducted for both distributions $P$ and $\dot{P}$ and yielded p-values of $0.24$ and $0.07$ respectively. Therefore, it is impossible to reject the null hypothesis that the simulated distributions are similar to the observed distribution.
@@
Fig 10 shows the good agreement between simulations and observations. Furthermore, we can compare now Fig. 11 and Fig. 7. It is obvious that Fig. 11 is better, because globally the colors are lighter which signifies the proportion of pulsars in each 2D bin are almost alike. In addition, there is not a pile up effect anymore next to the death line, which corresponds better to the data.
@@
The distribution of the distance to Earth is shown in Fig. 12. It has a slightly better similarity to the observations than the work of Dirson et al. (2022). Nevertheless the improvement is marginal compared to the improvements of $P$ and $\dot{P}$ histograms. However, we noticed that removing the effect of dispersion and scattering of the ISM on the radio pulse profile, we would obtain very similar distribution of distances between observations and simulations. Therefore, a refinement in the description of the effect of the ISM on the pulse profile might be necessary.
Indeed, Fig. 13 confirms that the positions of the pulsars in the X-Y plane of the Galaxy are slightly more scattered than before, but the improvement is marginal.
@@
Finally fig. 14 compares the Galactic latitudes from the previous study (see Fig. 9 of Dirson et al. (2022)), now the simulated distribution resembles more the observations, especially around all the latitudes that are different from 0°. Therefore, the Galactic potential helped tracking the pulsar trajectories more realistically.
@@ 3.2.2 Without a death line
Because the simulation where 1 million pulsars are simulated do not allow to obtain pulsars older than $4.1\times 10^{7}$ yr, we undertook a thorough examination of the usefulness of a death line in modeling the canonical pulsar population. Upon conducting simulations with and without a death line (see Table 5 for the parameter values used in the simulations), it emerged that the resultant number of detected pulsars and the corresponding plots exhibited remarkable similarities. In addition, when we conduct the KS test on $P$ and $\dot{P}$ between 2 simulations with and without death line, we get p-values of 0.71 and 0.16 respectively, which confirms the similarity between the two simulations. Consequently, our attention turned towards exploring the impact of lower birth rates to ascertain if they provide a more congruous fit for both spin period and spin period derivative distributions.
@@
Running simulations with a birth rate of $1/70$ yr-1 or a birth rate of $1/150$ yr-1 (as a consequence, increasing the oldest simulated pulsar), conspicuously revealed an insufficient number of detected pulsars in both Fig. 15 and Fig. 16. Intriguingly, a discernible paucity of pulsars congregating in the bottom right quadrant of the diagram, colloquially known as the pulsar graveyard, persisted even under these conditions, therefore regardless of the birth rates, thus casting doubt on the necessity of a death line in this configuration. Ultimately, our analysis underscores the pivotal importance of delineating the maximum age of simulated pulsars, as elucidated in Subsection 3.1, while the incorporation of a death line do not seem indispensable for attaining realistic results when simulating 1 million pulsars.
@@ 3.3 Spin-velocity misalignment effect
With the aim of observing the effect of the Galactic potential, we looked at the alignment of the rotation axis and the proper motion velocities of the pulsars. Results are taken from the simulation containing 1 million pulsars, see Table 5 for the parameters used. Noutsos et al. (2013) showed by studying a catalog of 58 pulsars that pulsars younger than 10 Myr keep their rotation axis and their velocities vector still aligned or anti-aligned, because they were aligned at birth according to Rankin (2007). Meanwhile, pulsars older than 10 Myr do not show this trend anymore. This progressive misalignment effect is due to the movement through the Galactic potential.
@@
Fig. 17 demonstrates that most pulsars have their rotation axis aligned with their velocities, but for spin-velocity angles between 25° and 180° the numbers of pulsars is considerably lower than around 0° and 180°. On the one hand, for pulsars younger than 10 Myr, shown in blue in Fig. 18, only a small fraction of pulsars have their spin-velocity angle greater than 25° and below 150°. On the other hand, for pulsars older than 10 Myr, shown in red in Fig. 18, their spin-velocity angle becomes greater than 25° or below 150°.
@@
Furthermore, in Fig. 19, we can confirm that, the older the pulsar is, the better is its chance to orbit more in the Galaxy, allowing it to deviate its spin-velocity angle from alignment (or anti-alignment). In addition, even if the pulsar is not that old, if it made many orbits in the Galaxy, its spin-velocity angle will be between 25° and 150°, which confirms the fact that its the Galactic potential (and not only age) which makes this spin-velocity angle increase (or decrease if its is anti-aligned at birth). In the contrary, there are pulsars which are old which did not made many orbits in the Galaxy, as a consequence they have their spin-velocity angle near 0° or 180°. These plots clearly shows the influence of the Galactic potential making the pulsars lose their alignment between their rotation axis and their velocities through time because they evolve in the Galactic potential.
@@ 3.4 Log normal vs normal distribution for spin period at birth
Figure 20: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation and the Gaussian distribution at birth for the spin period.
@@
Table 9: Number of pulsars detected with the death line implementation and the Gaussian distribution at birth for the spin period.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 7 0 5 2
¿28 275 9 138 128
Total 1642 1188 223 231
@@
In order to estimate the impact of a new distribution of spin period at birth, we compare the results with the previous distribution of spin period
$p(P_{0})=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(P_{0}-\bar{P})^{2}/(2\sigma_{p})^{2}}$ (34)
@@
The simulation in this subsection uses $\bar{P}=60$ ms and $\sigma_{p}=10$ ms as parameters of equation (34). The Galactic potential and the death line were both taken into account to compare the influence of the initial spin period distribution and 1 million pulsars were simulated, see Table 5 for the parameters used.
Table 9 highlights the fact that much more pulsars are detected above $\dot{E}>10^{28}$ W compared to the simulation with the Gaussian in log space for the spin period, see Table 7. It confirms the suspicion that the distribution of initial spin period influences the number of pulsars at higher spin down luminosity.
The spin period distribution in Fig. 20 shows some simulated pulsars with a low spin period, in excess compared to the observations. This artifact disappears when the Gaussian distribution in log space for the spin period is used. The log-normal distribution enables to remove those pulsars in excess and make the distribution of the spin period better fit the observations. The KS test was also conducted on both $P$ and $\dot{P}$ and we find p-values, once again, considerably below 0.05 for both quantity, therefore rejecting the null hypothesis. We conclude that the initial distribution for the spin period found by Igoshev et al. (2022) allows to better reproduce the canonical pulsars population.
3.5 Gamma-ray detection
Let us now focus on the $\gamma$-ray band of the pulsar population. First, we note that the total number of pulsars detected simultaneously in gamma-ray and in radio is significantly larger in our simulation compared to the observations, when using the Fermi/LAT instrument sensitivity. However, the simulated gamma-ray pulsars are in the expected area of Fig. 21, meaning the striped wind model allow to correctly reproduce these pulsars. The explanation of this higher number of detection would be that there are many unidentified sources in the Fourth Fermi-LAT Catalog of Gamma-ray sources (4FGL). Since many unidentified Galactic sources are listed in these catalogs, many gamma-ray pulsars might have been detected without being identified as such yet. Furthermore, we can compare our results with the PPS work of Pierbattista et al. (2012), where they simulated the young gamma-ray pulsars population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models. In summary their best model was the OPC model, which shows almost identical results as ours (the $P-\dot{P}$ shape and the number of detection), except the fact that we detect gamma-ray pulsars a little further away from Earth.

### v2

@@ 3.1.2 With the death line
Adding a death line decreases the number of pulsars detected in radio compared to the situation without a death line, see Table 7. The death line improves the fit as seen in Fig. 5 for the distribution of spin period and compared with Fig. 2, even though there is a slight excess of pulsars which have spin period between $10^{-0.5}$ s and $1$ s. In Fig. 6, these pulsars in excess lie at $\dot{P}$ between $10^{-18}$ and $10^{-17}$, where there is no data and which correspond to the oldest pulsars in the simulation. Therefore the oldest pulsars simulated here do not correspond to any real data. In addition, too many pulsars end up close to the death line, disagreeing with the observations. The death line creates a pile up effect not seen in the observations.
@@
Figure 7: Density plot of the $P-\dot{P}$ diagram in comparison with observations.
The goodness-of-fit is checked against a density plot of the $P-\dot{P}$ diagram as in Johnston & Karastergiou (2017) instead of putting points for individual pulsars. In order to construct such diagrams we binned the data evenly in a log scale in $P$ and $\dot{P}$ and count the number of pulsars in each bin. This leads to a 2D histogram to be compared with the observed density plot 2D histogram. Because the total number of simulated pulsars $N_{totsim}$ is different from the total number of observed pulsars $N_{totobs}$, we normalise the number of pulsars in each bin in order to compare proportions. We therefore introduce the quantity $R$ as follows
@@
With this definition the number $R$ remains in the interval $[-1,+1]$. $N_{sim}$ is the number of pulsar count in a bin of the simulation, $N_{obs}$ the number of pulsar detected in a bin of the observations. A good fit in a bin corresponds to a value of $R$ close to 00, meaning that the simulation results are close to the observations. Inspecting fig. 7, we note a stronger deviation from the observations at the boundaries of the density plot, confirming for instance the pile up close to the death line. The high values of $R$ at these boundaries originates from the weak number of pulsars counted in these bins leading to bad statistics and large variations in the proportions, a side effect of the density plot approach. The fact that the colors are very dark, shows that the proportion of pulsars in most of the bins are not good.
A Kolmogorov Smirnov (KS) test (Smirnov, 1948) was conducted on both $P$ and $\dot{P}$ which gave p-values substantially below 0.05 for both quantities, when no death line is implemented. While we obtained p-values of 0.05 for $P$ and well below 0.05 for $\dot{P}$ with a death line implementation. The null hypothesis that observations and simulations originate from the same distribution for $P$ can not be rejected (p-value $\geq 0.05$) only when a death line is considered, but is clearly rejected for $\dot{P}$ in both case.
@@ 3.2.1 With the death line
The parameters used for these runs are shown in Table 5. This new set of simulations is intended to verify if generating pulsars younger than $4.1\times 10^{7}$ yr is sufficient, since older pulsars are barely observed. The death line is also taken into account. Without any specifications, in all figures, with the exception of fig. 12, 13 and 14, simulations were performed while considering the ISM effect.
@@
Table 8 shows that a few less pulsars were detected in this simulation compared to the one with the death line and 10 millions pulsars (see Table 7). Moreover, inspecting the histograms for spin period and spin period derivative respectively in Fig. 8 and 9, we conclude that observations and simulations are very similar in both histograms. A KS test was conducted for both distributions $P$ and $\dot{P}$ and yielded p-values of $0.67$ and $0.17$ respectively. Therefore, it is impossible to reject the null hypothesis stating that the simulated distributions are similar to the observed distribution.
@@
Fig 10 shows the good agreement between simulations and observations. Furthermore, comparing Fig. 11 and Fig. 7, the former performs better because globally the values of $R$ are close to zero, highlighting the fact that the proportion of pulsars in each 2D bin are almost alike. In addition, the pile up effect close to the death line has been removed from the $P-\dot{P}$ diagram, which better fits to the data.
@@
The pulsar distance distribution to Earth is shown in Fig. 12. The improvement compared to Dirson et al. (2022) is marginal. Anyway, let us remind that the estimated distances of the observed pulsars are subject to large uncertainties because of the distances derived from the dispersion measure values (DM) are highly uncertain, while the distances for the simulated pulsars are correct by construction. As a consequence, the comparison between the observed and simulated distance distribution is not a reliable indicator to check the correctness of our model as the discrepancies are subject to observational uncertainties. Fortunately the $P-\dot{P}$ improved compared to the previous work. This could mean the observed pulsars are closer than expected. Indeed, Fig. 13 shows that the positions of the simulated pulsars in the X-Y plane of the Galaxy, including the ISM effect. They are closer to the Sun than the observed pulsars. Surprisingly, by removing the effect of dispersion and scattering of the ISM on the radio pulse profile, we obtain very similar distance distributions between observations and simulations, see Fig. 12. Therefore, a refinement of the ISM impact on radio pulse profiles might help getting a more realistic model.
@@
Finally fig. 14 compares the Galactic latitude distributions with the study of Dirson et al. (2022), their Fig. 9. In our work the simulated distribution better matches the observations, especially around all the latitudes away from 0°. Therefore, the Galactic potential helped tracking the pulsar trajectories more realistically. We notice also the ISM removes pulsars close to a latitude of 0°, because the electron density is highest in the galactic plane of the Milky Way.
@@ 3.2.2 Without a death line
Simulation with 1 million simulated pulsars does not allow to generate pulsars older than $4.1\times 10^{7}$ yr. We therefore undertook a thorough examination of the usefulness of a death line in modeling the canonical pulsar population. Upon conducting simulations with and without a death line (see Table 5 for the parameter values used in the simulations), we found that the resultant number of detected pulsars and the corresponding plots exhibited remarkable similarities. By conducting the KS test on $P$ and $\dot{P}$ between two simulations, one with and one without the death line, we obtained p-values of 0.20 and 0.16, respectively. Since both p-values are greater than 0.05, this indicates no significant difference between the two simulations, supporting their similarity. Consequently, our attention turned towards exploring the impact of lower birth rates to ascertain if they provide a more congruous fit for both spin period and spin period derivative distributions.
@@
Running simulations with a birth rate of $1/70$ yr-1 or a birth rate of $1/150$ yr-1 (as a consequence, increasing the oldest simulated pulsar), conspicuously revealed an insufficient number of detected pulsars in both Fig. 15 and Fig. 16. Intriguingly, a discernible paucity of pulsars congregating in the bottom right quadrant of the diagram, colloquially known as the pulsar graveyard, persisted even under these conditions, therefore regardless of the birth rates, thus casting doubt on the necessity of a death line in this configuration. Ultimately, our analysis underscores the pivotal importance of delineating the maximum real age of simulated pulsars, as elucidated in Subsection 3.1, while the incorporation of a death line do not seem indispensable for attaining realistic results when simulating 1 million pulsars.
@@ 3.3 Spin-velocity misalignment effect
With the aim of observing the effect of the Galactic potential, we looked at the alignment of the rotation axis and the proper motion velocities of the pulsars. Results are taken from the simulation containing 1 million pulsars, see Table 5 for the parameters used. Noutsos et al. (2013) showed by studying a catalogue of 58 pulsars that pulsars younger than 10 Myr keep their rotation axis and their velocities vector still aligned or anti-aligned, because they were aligned at birth according to Rankin (2007). Meanwhile, pulsars older than 10 Myr do not show this trend anymore. This progressive misalignment effect is due to the movement through the Galactic potential.
Fig. 17 demonstrates that most pulsars have their rotation axis aligned or counter-aligned with their velocities, but for spin-velocity angles between 30° and 150° the numbers of pulsars is considerably lower than around alignment 0° and counter-alignment 180°. On the one hand, for pulsars younger than 10 Myr, shown in blue in Fig. 18, only a small fraction of pulsars have their spin-velocity angle greater than 30° and below 150°. On the other hand, for pulsars older than 10 Myr, shown in red in Fig. 18, their spin-velocity angle becomes greater than 30° or smaller than 150°.
Furthermore, Fig. 19 confirms that, the older the pulsar is, the better its chance to perform several orbits in the Galaxy, allowing it to significantly deviate from spin-velocity alignment or anti-alignment. In addition, even for less old pulsars able to perform many orbits in the Galaxy, their spin-velocity angle will evolve between 30° and 150°. This supports the idea that the Galactic potential is responsible for the misalignment irrespective of the age of the pulsar. The opposite is also true, some old pulsars do not perform many orbits (but most of them are), and as a consequence they keep approximately their spin-velocity alignment or misalignment. The previous plots clearly highlight the influence of the Galactic potential making pulsars lose their alignment between their rotation axis and their proper motion velocities through time.
@@ 3.4 Log normal vs normal distribution for spin period at birth
Table 9: P-values and total number of pulsars detected ($N_{detection}$) obtained for simulations with different $\bar{P}$ for the Gaussian initial distribution of spin period.
$\bar{P}$ (ms) $\sigma_{p}$ (ms) p-value (P) p-value ($\dot{P}$) $N_{detection}$
60 30 $6.58\times 10^{-20}$ $3.76\times 10^{-4}$ 1512
80 30 $1.84\times 10^{-15}$ 0.051 1671
100 30 $2.34\times 10^{-10}$ 0.123 1721
129 30 $1.02\times 10^{-6}$ 0.050 1787
140 30 $1.83\times 10^{-5}$ 0.006 1816
@@
In order to estimate the impact of the spin period distribution at birth, we compare the results of the log-normal law with the previously used normal law in Dirson et al. (2022)
$p(P_{0})=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(P_{0}-\bar{P})^{2}/(2\sigma_{p})^{2}}\ .$ (36)
@@
The simulation in this subsection uses the parameters listed in the two first columns of Table 9 and to be put into equation (36). The Galactic potential and the death line were both taken into account to compare the influence of the initial spin period distribution on the 1 million pulsars simulated, see Table 5 for the parameters used. In Table 9, only results for $\sigma_{p}=30$ ms are shown because other values have been tried out, namely 10 ms, 20 ms and 40 ms but gave no higher p-values than for $\sigma_{p}=30$ ms. Moreover, we chose $\bar{P}$ from 60 ms to 140 ms because it is the range usually used in other work (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Dirson et al., 2022).
Table 9 highlights the fact that for $\bar{P}$ above 60 ms and below 140 ms, an acceptable p-value $\geq 0.05$ is reached for the $\dot{P}$ distribution. However no matter the value of $\bar{P}$, a satisfying p-value for the $P$ distribution is not reached in contrary to the log-normal distribution at birth, for which satisfying p-value for both distributions are obtained. We conclude that the initial distribution for the spin period found by Igoshev et al. (2022) allows to better reproduce the canonical pulsars population.
3.5 $\gamma$-ray detection
Let us now focus on the $\gamma$-ray pulsar population. First, we note that the total number of pulsars detected is significantly larger in our sample (307 detections) compared to the observations (173 detections), when using the Fermi/LAT instrument sensitivity. However, the simulated $\gamma$-ray only and radio loud $\gamma$-ray pulsars are in the expected area of Fig. 20, furthermore, the KS test gives a p-value of 0.52 and $\approx 0$ for $\dot{P}$ and $P$ respectively. The null hypothesis can not be rejected for $\dot{P}$ but it is rejected for $P$. However, when the KS test is conducted on the $\gamma$-ray population only we get p-values of 0.03 and 0.14 for $\dot{P}$ and $P$ respectively, therefore here the null hypothesis is rejected for $\dot{P}$ but not for $P$. While for the radio loud $\gamma$-ray population p-values of 0.11 and $\approx 0$ are obtained for $\dot{P}$ and $P$ respectively. Thus, we thought the low p-value obtained for $P$ for both population was probably because of the 7 high spin period radio loud $\gamma$-ray pulsars seen on Fig. 20, but even when conducting the KS test without these pulsars gave a p-value $\geq 0.05$ for $\dot{P}$ and well below $0.05$ for $P$. In addition, we also compare the $\gamma$-ray flux distribution of the simulated and observed populations. In Fig.21 we notice that both distributions peak at the same flux level around $10^{-13.7}$ W.m-2. Globally, the number of pulsars in the simulation with a flux greater than $10^{-13.4}$ W.m-2 is very close to the observations. However, between $10^{-14.4}$ W.m-2 and $10^{-14}$ W.m-2, there are much more pulsars in the simulation. This bi-modality in the simulated $\gamma$ray flux distribution is an artefact introduced by our abrupt condition on the threshold $F_{\rm min}$ depending on the radio detection of a pulsar and on a sufficiently low galactic latitude. However if we smoothly change $F_{\rm min}$ from $4\times 10^{-15}$ to $16\times 10^{-15}$ (the two $F_{\rm min}$ values used in this work), the bi-modality would almost disappear. The higher number of detected pulsars in the simulations could be related to the unidentified sources in the Fourth Fermi-LAT catalogue of $\gamma$-ray sources (4FGL). Since many unidentified Galactic sources are listed in this catalogue, many $\gamma$-ray pulsars might have been observed without being identified as such yet. Nonetheless, assuming that every unidentified $\gamma$-ray pulsar in this catalogue will be recognized as such by a future more sensitive instrument and that a significant difference between the number of simulated and observed pulsars subsists, this could hint to some shortcomings faced by the striped wind model. Let us mention some of these possible flaws of geometrical or physical nature. First the $\gamma$-ray beam opening angle could be smaller than the one used due to some plasma recollimation effect within the striped wind. This would reduce the number of detected pulsars. Second the luminosity function could also deviate from the prescription adopted in (32). Another dependence on $B$ and $P$ or equivalently on $P$ and $\dot{P}$ could be introduced, maybe adding a third parameter like the cut-off energy to better describe the $\gamma$-ray luminosity. Furthermore, we compare our results with the PPS work of Pierbattista et al. (2012), where they simulated the young $\gamma$-ray pulsar population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models. To summarize their findings, their best model is the OPC model, showing almost identical results as ours, $P-\dot{P}$ shape and number of detection, except that we detect $\gamma$-ray pulsars a little farther away from Earth.